OpenStudy (yumeybitaa):

Evaluate 62 ÷ (3 + 9). 3 4 12 21

OpenStudy (anonymous):

5.17

OpenStudy (shadowlegendx.):

@sheelo_mughal it is against the CoC to give out plain answers. This site is designed for helping users, not giving them plain answers and they learn nothing

OpenStudy (shadowlegendx.):

@Yumeybitaa are you there?

OpenStudy (shadowlegendx.):

As I told you the other day, PEMDAS is key to this type of equation

OpenStudy (shadowlegendx.):

Notice how (3 + 9) is in parentheses, that means we must do this section first. I believe you would know what 3 + 9 is so I will lay the next step out for you. \[62 \div 12 = ? \]

OpenStudy (crashonce):

62/(3+9) =62/12 =5.17 approx

OpenStudy (shadowlegendx.):

Crash I appreciate you adding your help to this question but usually I like to have the asker solve the equation, not other people :)

OpenStudy (shadowlegendx.):

It helps them learn

OpenStudy (yumeybitaa):

i still dont get it

OpenStudy (crashonce):

sorry ShadowLegend im new lol

OpenStudy (yumeybitaa):

help please ?

OpenStudy (crashonce):

basically simplify whats in the brackets (3+9)

OpenStudy (crashonce):

then that is 12. now divide 62 by 12

OpenStudy (shadowlegendx.):

Welcome to OpenStudy @CrashOnce :)

OpenStudy (yumeybitaa):

5.16666667

OpenStudy (shadowlegendx.):

This isnt my normal account, my main one is @ShadowLegendX c;

OpenStudy (yumeybitaa):

??

OpenStudy (crashonce):

yes your answer is correct @Yumeybitaa